# Fermion masses and mixings in supersymmetric SO(10)

### v1

@@ 2.1 Quasi Yukawa Unification
where the latter ratio is independent of $\tan\beta$. This suggests that an additional contribution to the third-generation fermion masses is necessary to account for the $b$$\tau$ splitting. We propose to achieve this via the following higher-dimensional operator:
@@
where $\Lambda$ represents a cutoff scale (we will later provide an estimate for this $\Lambda$). The non-renormalizable term in Eq. (2.5) can provide a relative Clebsch-Gordan (CG) coefficient of $(-3)^{2}$ for the tau-lepton (for a systemic analysis of CG factors in SO(10) setup, see Refs. [53, 54]). Interestingly, this somewhat large CG factor of $+9$ can overcome the $M_{\mathrm{GUT}}^{2}/\Lambda^{2}$ suppression in the above operator and provide the desired tau-bottom Yukawa coupling ratio as quoted above. For the operator in Eq. (2.5), we may assume that it appears from integrating out vectorlike $16+\overline{16}$ fermions, which fixes the types of contractions we allow.
The desired CG coefficient can be obtained if the adjoint Higgs $45_{H}$ gets a vacuum expectation value (VEV) in the $B-L$ direction, namely,
@@
[…]
$16_{H}+\overline{16}_{H}]{c,\overline{c}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}.$ (2.8)
@@ 2.2 Structure of the 2–3 Sector
This term provides a relative CG coefficient of $-3$ for the charged leptons since $\langle 45_{H}\rangle\propto$ B-L (anti-symmetric due to the VEV structure $\langle 45_{H}\rangle\propto$ B-L), which plays a crucial role in obtaining the above-mentioned desired ratio for the second-generation Yukawa couplings. For the operator in Eq. (2.12), we may assume that it appears from integrating out vectorlike $16+\overline{16}$ fermions, which fixes the types of contractions we allow.
@@ 3 Numerical Analysis
$y_{t}\sin\beta$ $0.5721$ $0.5679$ $-0.144$
$y_{d}\cos\beta/10^{-5}$ $0.6866$ $0.5114$ $-\underline{1.27}$
@@
$\delta^{\circ}_{\mathrm{PMNS}}$ 16.62 -
$\chi^{2}$ 8.8
@@
In this section we provide a full numerical analysis of our proposed fermion masses as given in Eqs. (2.24), (2.25), and (2.30). These mass matrices are given at the GUT scale. Therefore, the experimentally measured values of the charged fermion masses and mixings to be fitted need to be taken at the GUT scale. For this purpose, we use the input values of the GUT scale ($M_{\mathrm{GUT}}=2\times 10^{16}$ GeV, as suggested by gauge coupling unification within the MSSM setup) as given in Ref. [52] corresponding to a scenario with zero threshold corrections from supersymmetric partners (moreover, for definiteness, we fix the supersymmetry breaking scale at 3 TeV). For observables with uncertainties $<5\%$, we have allowed a maximum error of $5\%$. Note, however, that threshold corrections from supersymmetric partners may, in some cases, introduce even larger deviations from these values.
@@
It is worth nothing here that the value of $\Lambda$ we have estimated in Eq. (3.8) is close to $M_{\mathrm{P}}/\sqrt{N}$, where $N$ denotes the total number of degrees of freedom in this model and $M_{\mathrm{P}}=2.44\times 10^{18}$ GeV is the reduced Planck mass. In our theory with $10_{H},45_{H},16_{H},\overline{16}_{H}$, and three generations of chiral superfields $16_{i}$, we have $N=540$ propagating particle species, which results in $\Lambda\simeq M_{\mathrm{P}}/\sqrt{N}\simeq 1.05\times 10^{17}$ GeV, consistent with Eq. (3.8). For a discussion of why the ultraviolet cutoff scale is $M_{\mathrm{P}}/\sqrt{N}$, see Refs. [64, 65].
@@
$y_{t}\sin\beta$ $0.5069$ $0.4992$ $-0.301$
$y_{d}\cos\beta/10^{-5}$ $0.4971$ $0.3839$ $-\underline{1.138}$
$y_{s}\cos\beta/10^{-4}$ $0.9838$ $0.9657$ $-0.367$
@@
$\delta^{\circ}_{\mathrm{PMNS}}$ 37.3 -
$\chi^{2}$ 17.14
@@ 4 Metastable/Quasistable Strings in SO(10)
In our discussion so far we have made a reasonable assumption that the Higgs multiplets $45_{H}$ and $16_{H}$, $\overline{16}_{H}$ acquire the same GUT scale VEVs. For this case, the SO(10) symmetry breaking only produces a superheavy magnetic monopole that carries a single unit ($2\pi/e$) of Dirac magnetic charge as well as some color magnetic charge [66]. In order to implement a metastable string scenario in SO(10) which entails a cosmic string that can be broken through quantum tunneling of a monopole-antimonopole pair, we assume the presence of an additional Higgs 45-plet (which we label by $45^{\prime}_{H}$) such that the SO(10) symmetry breaking proceeds as follows:
$SO(10)$ ${45_{H}}SU(3)_{c}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$ (4.1)
${45^{\prime}_{H}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{R}\times U(1)_{B-L}$ (4.2)
${16_{H}+\overline{16}_{H}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}\;.$ (4.3)
@@ 5 Conclusions
We have described how the observed fermion masses and mixings can be accommodated in a supersymmetric SO(10) model based on a minimal Higgs sector. In addition to the renormalizable Yukawa couplings, an essential role is played by a suitable choice of non-renormalizable Yukawa couplings. A numerical analysis shows that the model has a preference for a solution with large $\tan\beta\sim 58$, which is compatible with the postulate of third family quasi-Yukawa unification. For our model this corresponds to the asymptotic relation $y_{t}:y_{b}:y_{\tau}=1+\zeta:1+\zeta:1-9\zeta$, where the coefficient 9 is a group-theoretic factor, and $\zeta$ is estimated from our fits to be close to 0.03. We also consider low $\tan\beta$ solutions and present one with $\tan\beta$= 10. The three right handed neutrino masses are estimated to be $10^{9}$, $8\cdot 10^{12}$, and $9\cdot 10^{12}$ GeV. Finally, we briefly discuss how the metastable and quasistable cosmic string scenarios can be realized in this class of realistic SO(10) models.

### v2

@@ 1 Introduction
In summary, the novel direction pursued in our present work includes the proposal to incorporate a specific non-renormalizable operator that introduces a group-theoretical factor of ‘9’ in the tau-lepton sector, thereby enabling the realization of third-generation quasi-Yukawa unification in the form $y_{t}:y_{b}:y_{\tau}=1+\zeta:1+\zeta:1-9\zeta$ (with $\zeta\simeq 0.03$). Our proposal provides a group-theoretical origin for achieving quasi-Yukawa unification for the third generation in supersymmetric scenarios where threshold corrections are negligible. To the best of our knowledge, this is the first demonstration within the SO(10) framework of a consistent fit to the full three-generation fermion masses and mixings—including both charged fermions and neutrinos—that predicts third-generation quasi-Yukawa unification, fully compatible with the current experimental data.
@@ 2.1 Quasi Yukawa Unification
[…]
@@
where $\Lambda$ represents a cutoff scale (we will later provide an estimate for this $\Lambda$). The non-renormalizable term in Eq. (2.5) can provide a relative Clebsch-Gordan (CG) coefficient of $(-3)^{2}$ for the tau-lepton (for a systemic analysis of CG factors in SO(10) setup, see Refs. [53, 54]). Interestingly, this somewhat large CG factor of +9+9 can overcome the $M_{\mathrm{GUT}}^{2}/\Lambda^{2}$ suppression in the above operator and provide the desired tau-bottom Yukawa coupling ratio as quoted above. For the operator in Eq. (2.5), we may assume that it appears from integrating out vectorlike $16+\overline{16}$ fermions, which fixes the types of contractions we allow. Note that a lower dimensional operator, namely $16_{3}16_{3}10_{H}45_{H}$, cannot break the equality $y_{b}=y_{\tau}$, since $10\times 45\supset 10+120$, with the effective $120_{H}$ not coupling to $16_{3}16_{3}$ owing to flavor antisymmetry.
The desired CG coefficient can be obtained if the adjoint Higgs $45_{H}$ gets a vacuum expectation value (VEV) in the BLB-L direction, namely,
@@
$SO(10)$ $\xrightarrow[45_{H}]{a}SU(3)_{c}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$ (2.7)
$\xrightarrow[16_{H}+\overline{16}_{H}]{c,\overline{c}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}.$ (2.8)
@@ 2.2 Structure of the 2–3 Sector
This term provides a relative CG coefficient of 3-3 for the charged leptons since $\langle 45_{H}\rangle\propto$ B-L (anti-symmetric due to the VEV structure $\langle 45_{H}\rangle\propto$ B-L), which plays a crucial role in obtaining the above-mentioned desired ratio for the second-generation Yukawa couplings. For the operator in Eq. (2.12), we may assume that it appears from integrating out vectorlike $16+\overline{16}$ fermions, which fixes the types of contractions we allow.
@@ 3 Numerical Analysis
$y_{t}\sin\beta$ $0.5721$ $0.5679$ 0.144-0.144
$y_{d}\cos\beta/10^{-5}$ $0.6866$ $0.5114$ $\underline{1.27}$
@@
$\delta^{\circ}_{\mathrm{PMNS}}$ - 16.62 -
$\chi^{2}$ - - 8.8
@@
In this section we provide a full numerical analysis of our proposed fermion masses as given in Eqs. (2.24), (2.25), and (2.30). These mass matrices are given at the GUT scale. Therefore, the experimentally measured values of the charged fermion masses and mixings to be fitted need to be taken at the GUT scale. For this purpose, we use the input values of the GUT scale ($M_{\mathrm{GUT}}=2\times 10^{16}$ GeV, as suggested by gauge coupling unification within the MSSM setup) as given in Ref. [52] corresponding to a scenario with zero threshold corrections from supersymmetric partners (moreover, for definiteness, we fix the supersymmetry breaking scale at 3 TeV). For observables with uncertainties $5\%$, we have allowed a maximum error of $5\%$. Note, however, that threshold corrections from supersymmetric partners may, in some cases, introduce even larger deviations from these values.
@@
It is worth emphasizing that the most minimal and predictive SUSY SO(10) frameworks—those that do not incorporate SUSY threshold corrections—are characterized by a total of 19 input parameters. This omission is well-motivated, as threshold corrections depend sensitively on the superpartner mass spectrum, which is theoretically unconstrained and introduces additional free parameters, thereby reducing the overall predictivity of the model. A recent example of such a minimal setup can be found in Ref. [41], where a complete and successful fit to the fermion mass spectrum yields a total $\chi^{2}\sim 4$. As demonstrated in our work, we also adopt a similarly minimal parameter set consisting of 19 inputs (14 magnitudes and 5 phases), and obtain a comparable fit with $\chi^{2}\sim 8$ (see Table 1). While alternative approaches in the literature do include SUSY threshold corrections, they typically require additional free parameters to achieve a successful fit, making them less constrained and less predictive in comparison. It is straightforward to include SUSY threshold corrections in our fitting procedure by enlarging the parameter space of the theory, which would likely lead to a further reduction in the total $\chi^{2}$.
@@
It is worth nothing here that the value of $\Lambda$ we have estimated in Eq. (3.8) is close to $M_{\mathrm{P}}/\sqrt{N}$, where $N$ denotes the total number of degrees of freedom in this model and $M_{\mathrm{P}}=2.44\times 10^{18}$ GeV is the reduced Planck mass. In our theory with $10_{H},45_{H},16_{H},\overline{16}_{H}$, and three generations of chiral superfields $16_{i}$, we have $N=540$ propagating particle species, which results in $\Lambda\simeq M_{\mathrm{P}}/\sqrt{N}\simeq 1.05\times 10^{17}$ GeV, consistent with Eq. (3.8). For a discussion of why the ultraviolet cutoff scale is $M_{\mathrm{P}}/\sqrt{N}$, see Refs. [64, 65]. Given a lower Planck scale, there could be significant smearing of the gauge couplings near the unification scale from operators of the type $G_{\mu\nu}G^{\mu\nu}45^{2}_{H}/\Lambda^{2}$ (note that a lower dimensional operator of the form $G_{\mu\nu}G^{\mu\nu}45_{H}/\Lambda$ vanishes due to the antisymmetric property of $45_{H}$). This effect could, in general, help with the unification of gauge couplings in the one-step breaking of the GUT symmetry to MSSM, because the higher-dimensional operators are expected to generate states that may lie slightly below the GUT scale, which could otherwise spoil successful coupling unification (for details, see Ref. [14]).
@@
$y_{t}\sin\beta$ $0.5069$ $0.4992$ 0.301-0.301
$y_{d}\cos\beta/10^{-5}$ $0.4971$ $0.3839$ $\underline{1.138}$
$y_{s}\cos\beta/10^{-4}$ $0.9838$ $0.9657$ 0.367-0.367
@@
$\delta^{\circ}_{\mathrm{PMNS}}$ - 37.3 -
$\chi^{2}$ - - 17.14
@@ 4 Metastable/Quasistable Strings in SO(10)
In our discussion so far we have made a reasonable assumption that the Higgs multiplets $45_{H}$ and $16_{H}$, $\overline{16}_{H}$ acquire the same GUT scale VEVs. For this case, the SO(10) symmetry breaking only produces a superheavy magnetic monopole that carries a single unit ($2\pi/e$) of Dirac magnetic charge as well as some color magnetic charge [66]. In order to implement a metastable string scenario in SO(10) which entails a cosmic string that can be broken through quantum tunneling of a monopole-antimonopole pair, we assume the presence of an additional Higgs 45-plet (which we label as $45^{\prime}_{H}$ and assume that it does not participate in Yukawa interactions; consequently, the previously established analysis of the charged fermion spectrum remains unaffected) such that the SO(10) symmetry breaking proceeds as follows:
$SO(10)$ $\xrightarrow[]{45_{H}}SU(3)_{c}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$ (4.1)
$\xrightarrow[]{45^{\prime}_{H}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{R}\times U(1)_{B-L}$ (4.2)
$\xrightarrow[]{16_{H}+\overline{16}_{H}}SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}\;.$ (4.3)
@@ 5 Conclusions
We have described how the observed fermion masses and mixings can be accommodated in a supersymmetric SO(10) model based on a minimal Higgs sector. In addition to the renormalizable Yukawa couplings, an essential role is played by a suitable choice of non-renormalizable Yukawa couplings. A numerical analysis shows that the model has a preference for a solution with large $\tan\beta\sim 58$, which is compatible with the postulate of third family quasi-Yukawa unification. For our model this corresponds to the asymptotic relation $y_{t}:y_{b}:y_{\tau}=1+\zeta:1+\zeta:1-9\zeta$, where the coefficient 9 is a group-theoretic factor, and $\zeta$ is estimated from our fits to be close to 0.03. The novel group-theoretical factor proposed in this work is essential for achieving compatibility with current fermion mass data in supersymmetric scenarios where threshold corrections are negligible. We also consider low $\tan\beta$ solutions and present one with $\tan\beta$= 10. The three right handed neutrino masses are estimated to be $10^{9}$, $8\cdot 10^{12}$, and $9\cdot 10^{12}$ GeV. Finally, we briefly discuss how the metastable and quasistable cosmic string scenarios can be realized in this class of realistic SO(10) models.